OpenStudy (anonymous):

is there any equation represent a line in 3-dimension?

OpenStudy (nottim):

not sure, but parametrics?

OpenStudy (anonymous):

Only parametric... \[\frac{x-x_1}{a}=\frac{y-y_1}{b}=\frac{z-z_1}{c}=t\]

OpenStudy (anonymous):

ok, thanks

OpenStudy (nottim):

aww, no medal for me. oh well. but parametrics are a form of another equation. might wanna look that up.

OpenStudy (anonymous):

i want to if possible but the medal system has been limited to for 1 person only.

OpenStudy (anonymous):

x=x0+at y=y0+bt z=z0+ct where t is a parameter What is the parameter mean?

OpenStudy (nottim):

i dunno. try wolframalpha.com

OpenStudy (nottim):

or use purplemath or calculushelp.com

OpenStudy (nottim):

u good?

OpenStudy (anonymous):

still learning :), thanks for the links

OpenStudy (nottim):

ok bye! goood luck eh.

OpenStudy (anonymous):

Yes, it can only be done parametrically. If you write something such as \(z=f(x,y)\), you're going to get a surface...no way to get a line out of that.

OpenStudy (nottim):

Like what about Vectors: {x,y,z]=[X0,y0,z0)+k(n1,m2,m3}

OpenStudy (nottim):

not totally sure tho. i failed that class

OpenStudy (anonymous):

That's exactly right... that's an example of a parametrization with respect to \(k\). If you go component-by-component and solve for \(k\), you will get the suggestion above.

OpenStudy (nottim):

heh...didnt actuall understand that.

OpenStudy (anonymous):

Simple answer: You are correct. ;-)

OpenStudy (nottim):

Wonderful!